OpenStudy (anonymous):

In a parking lot of 240 red and blue cars, the ratio of red cars to blue cars is 3:5. How many red cars are in the parking lot?

OpenStudy (mathmale):

I'd suggest you set up an equation involving two ratios. The first ratio would be that of the number of red cars to the number of blue cars; that's 3:5. We get on the left side of our equation\[\frac{ 3 }{ 5 }\] ... where that 3 refers to the number of red cards and that 5 refers to the number of blue cars.

OpenStudy (mathmale):

On the right side, you'd need another ratio, this one involving r, the number of red cars (out of 240 total) in the lot. Note that we can eliminate b, the number of blue cars, by taking advantage of the fact that r + b = 240: r = 240-b. Then \[\frac{ r }{ b }~becomes~\frac{ r }{ 240-r },\] and this ratio is equal to that on the left side:\[\frac{ 3 }{ 5 }=\frac{ r }{ 240-r }\] All you have to do now is to solve this for r. Ask questions about any aspect of this approach that is not clear for you.